OpenStudy (anonymous):

Multiply or Divide. State any restrictions on the variables. (x + 2/ x - 1)/(x + 4/ x^2 + 4x - 5)

hero (hero):

Hint: \(\large\frac{x+2}{x-1} \div \frac{x+4}{x^2 +4x - 5}\) \( \large= \frac{x+2}{x-1} \times \frac{x^2 + 4x - 5}{x+4}\)

hero (hero):

Also... \(x^2 + 4x - 5 = (x - 1)(x + 5)\)

hero (hero):

@living4him, Did you figure out what the answer is?

OpenStudy (anonymous):

Yeah I did, thanks!

hero (hero):

Well, let us know what it is :D

hero (hero):

@living4him

OpenStudy (anonymous):

ok its \[(x + 2)(x + 5)/x + 4, x \neq 1, -4\]

hero (hero):

There's a small issue with the restrictions for this.

hero (hero):

It is possible for -5 to be a restriction

OpenStudy (anonymous):

oh, well it says I got the answer right

hero (hero):

Awesome :D